### v1

# Keyword-Aware ASR Error Augmentation for Robust Dialogue State Tracking

Abstract
Dialogue State Tracking (DST) is a key part of task-oriented dialogue systems, identifying important information in conversations. However, its accuracy drops significantly in spoken dialogue environments due to named entity errors from Automatic Speech Recognition (ASR) systems. We introduce a simple yet effective data augmentation method that targets those entities to improve the robustness of DST model. Our novel method can control the placement of errors using keyword-highlighted prompts while introducing phonetically similar errors. As a result, our method generated sufficient error patterns on keywords, leading to improved accuracy in noised and low-accuracy ASR environments.
1 Introduction
Figure 1: Illustration of EPA process.
Task-oriented dialogue systems (TODs) assist users in achieving specific objectives through conversations and are used in various sectors, including customer service and hotel reservations. A crucial component of these systems is Dialogue State Tracking (DST), which extracts vital information from conversations in a slot-value format (e.g., hotel-name: Claire Hotel). This information is essential for querying databases and generating responses Young et al. (2013).
However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR). Notably, Soltau et al. (2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments [^1] Joint goal accuracy result of base size model.. This decline is primarily due to ASR errors, which frequently misrecognize named entities—a key target in DST Nechaev et al. (2021).
To address ASR inaccuracies, data augmentation has emerged as a viable, cost-efficient strategy. Existing text augmentation methods, such as word swapping Wei and Zou (2019) and back translation Sennrich et al. (2015), do not maintain audio similarity with the original text, leading to discrepancies with ASR error patterns. To bridge this gap, Sharma et al. (2020) and Jacqmin et al. (2023) synthesized audio from text with text-to-speech (TTS) model Shen et al. (2018) and processed it through ASR, while Hrinchuk et al. (2020) and Zhang et al. (2021) employed translation model structure to introduce ASR-like errors directly into texts.
Despite these advancements, prior methods often fail to provide sufficient error patterns for DST model training. Accurately identifying key terms is vital for DST performance; thus, models need to be trained on a broad spectrum of ASR-errored keywords. Unfortunately, many current strategies do not ensure that errors are positioned within critical keywords, often generating trivial examples by altering non-essential words such as random words Wei and Zou (2019) or sentence structure Sennrich et al. (2015). This oversight results in sub-optimal DST performance against ASR errors.
To address these limitations, we introduce Error Positioning Augmentation (EPA), a straightforward yet effective method that ensures sufficient errors in keywords. Our method leverages large language models (LLMs) Ouyang et al. (2022); Touvron et al. (2023); Zhang et al. (2022), which have demonstrated impressive capabilities in semantic augmentation Whitehouse et al. (2023); Sahu et al. (2023) and precise text generation control Sun et al. (2023); Liang et al. (2024). Despite their strengths, LLMs’ potential for phonetic augmentation remains largely unexplored.
In our method, we utilize in-context Brown et al. (2020) learning to guide the augmentation process. By providing examples that highlight the key parts, we enable the model to position errors in keywords. Surprisingly, without collecting extensive user voice datasets, a manually created small text dataset (fewer than 10 samples) can generate a wide variety of ASR-errored keywords. This simplifies the error generation process. Additionally, before keyword-focused augmentation, we augmented the text to incorporate utterance-level overall ASR errors, introducing diverse errors from open-source audio datasets. This two-step approach effectively harnesses LLMs for both overall utterance-level and targeted keyword-level phonetic augmentation.
In our experiments, EPA achieved a diversity in targeted keywords (95.4%) and high variance in overall text (1.81x) while maintaining audio similarity (91.5%). During practical DST evaluations, EPA significantly enhanced the robustness of DST models, increasing accuracy from 45.76% to 51.12%. Our analysis indicates that these improvements are primarily due to keyword augmentation, which significantly reduces errors in ASR-errored values.
2 Method
2.1 Notation
Before detailing each step, we first clarify the notation. Dialogue context from turn 1 to $t$ is denoted as $D_{t}$=$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance. DST model predicts the dialogue state (also called belief state) $B_{t}$ given $D_{t}$. $B_{t}$ is composed with slot $sl$ and value $v$ pairs, denoted as $B_{t}$ = $\{({sl}_{1},v_{1}),...,({sl}_{J},v_{J})\}$ , where ${sl}_{j}$ and $v_{j}$ is $j$-th slot name and value. $J$ is the total number of slots.
Idx Method Examples
1 Original I’d like to find a vegetarian restaurant, if possible.
+EPA I’d hike to find a veggie tarian restroom, if possible.
2 Original I am going to auburn.
+EPA I am flowing to auburng.
3 Original Tuesday, going to bailey’s crossroads please.
+EPA Tuesday, going to baley’s crossroads, peas.
4 Original Hi! Could you find me a train to loris on thursday?
+EPA Oh! Could you find me a trai to lorri on thursdae?
5 Original Ashby is my destination.
+EPA Ashy’s my desity.
Table 1: Examples of ASR errors from EPA. We have highlighted utterance level overall errors in yellow and keyword-specific errors in blue.
2.2 Step 1: ASR Error for Overall Utterance
In this step, we augmented the overall utterance by introducing general ASR errors. We began by constructing example sets for in-context learning, utilizing an open-source audio dataset Ardila et al. (2020). From this dataset, we randomly selected 300 hours of audio along with their corresponding gold transcripts ($g$) and transcribed the audio using an off-the-shelf ASR model (e.g., Whisper-base Radford et al. (2022)) to obtain the erroneous transcriptions ($e$). We denote this example dataset as $DB=\{(g_{1},e_{1}),\ldots,(g_{I},e_{I})\}$.
Next, we augment the overall utterance level of $u$ by guiding the LLM using examples retrieved from $DB$ based on audio similarity (see Figure  1, Step 1). To calculate audio similarity, we converted the characters to phonemes using the International Phonetic Alphabet (IPA) 222 https://pypi.org/project/eng-to-ipa/ and then computed the similarity between the phonemized $u$ and $DB$. After selecting the $k$ examples, we concatenated the instructions, examples, and $u$ into a single prompt and provided it to the LLM. This process resulted in the overall ASR-errored user utterance $\dot{u}$. Concretely, $\dot{u}$ can be obtained by
$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t}).$ (1)
2.3 Step 2: ASR Error for Keywords
Method Features Low-acc ASR Noised Aud. Paraphrased High-acc ASR
Aud. Utt-aug Key-aug LLM JGA N-Acc JGA N-Acc JGA N-Acc JGA N-Acc
Baseline - - - - 29.88 45.76 29.70 46.77 28.92 48.79 34.87 52.07
AEDA Karimi et al. (2021) - - - 29.90 46.46 29.74 47.48 29.12 48.86 34.94 52.32
EDA Wei and Zou (2019) - - - 29.22 47.65 28.70 49.51 28.08 49.99 33.68 53.78
BT Sennrich et al. (2015) - - - 31.69 49.17 31.26 50.98 29.90 51.73 36.27 54.81
TTS-ASR - - 29.94 46.34 29.99 47.37 29.08 48.88 35.07 52.03
ASR-translation - - 30.40 47.65 30.14 48.45 29.54 50.38 35.25 53.66
[…]
   w/o Keyword Aug - 31.43 49.63 31.51 50.56 30.41 52.02 36.34 54.57
EPA (Llama 7B) 31.54 51.12 31.55 52.27 30.10 53.55 36.22 55.49
   w/o Keyword Aug - 31.12 50.33 31.44 52.07 30.01 53.49 35.70 54.90
EPA (GPT3.5) 32.39 51.12 32.24 52.70 30.95 53.34 36.61 55.87
   w/o Keyword Aug - 31.31 50.67 31.13 52.29 30.06 52.85 35.40 55.80
Olisia Jacqmin et al. (2023) - - - - 30.17 46.25 30.43 48.07 29.13 49.21 36.1 52.58
Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments. All results were averaged over three seeds for better consistency.
While Step 1 introduces general ASR errors into ${u}$, it does not ensure that these errors enhance the diversity of keywords. To create a more effectivetraining dataset, we generated keyword-focused ASR errors (see Figure 1, Step 2).Here, we manually created several examples from the original DST dataset by explicitly highlighting the keywords in the clean text using the <hl> tag ($\tilde{g}$) and concatenated them with the ASR-errored versions ($\tilde{e}$). We set the $v$ in turn $t$ as keywords. We applied the same highlighting to the inference target $\dot{u}$ and then concatenated the instructions, examples and highlighted $\dot{u}$ into a single prompt, as in Step 1. We perform the in-context learning with this prompt to generate the augmented utterance $\ddot{u}$, containing overall ASR and keyword-specific errors. Explicitly, we get $\ddot{u}$ by
$\ddot{u_{t}}=LLM(Inst_{2}\oplus(\tilde{g}_{0},\tilde{e}_{0})\cdots(\tilde{g}_{k},\tilde{e}_{k})\oplus\dot{u}_{t}).$ (2)
We set the example numbers as 3 in both Step1 and Step2.
After generating the ASR-errored examples, we trained a DST model using both the EPA-generated data and the original seed data. Specifically, we fine-tuned T5-base Roberts et al. (2019) models for DST tasks, with detailed training procedures outlined in the Appendix B.
3 Augmentation Results of EPA
We first assessed the quality of ASR errors simulated by the EPA method to guarantee they closely resembled genuine ASR errors. Table 1 demonstrates that with overall utterance and keyword error augmentation, EPA effectively generates sentences that align with our objectives: maintaining phonetic similarity to the original text while confirming errors in keywords.
4 Experiments
4.1 Experimental Setup
Dataset. The DSTC11 dataset Soltau et al. (2022), an audio version of MultiWOZ 2.1 Eric et al. (2019), comprises 8,000 dialogues for training, 1,000 for validation, and 1,000 for testing. To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by word error rate (WER) [^3] Based on LibriSpeech.test-clean Radford et al. (2022). and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model (WER < 0.03) enviroment.
Metrics. For overall performance evaluation, we used joint goal accuracy (JGA), which requires all slot-value pairs to match the gold label. We also reported named entity accuracy (N-acc), the average accuracy across named entity slots.
LLMs. we primarily used GPT-3.5 Ouyang et al. (2022). Additionally, as this is the first attempt to augment pronunciation-related errors with LLMs, we also included Llama2-7B Touvron et al. (2023) and OPT-6.7B Zhang et al. (2022).
Compared methods. We compared our method with two established approaches: text-based augmentations, AEDA Karimi et al. (2021), EDA Wei and Zou (2019), and Back Translation (BT) Sennrich et al. (2015), and audio-aware augmentation methods, using synthesized audio (TTS-ASR) and translation model structure (ASR-translation). Lastly, we included Olisia Jacqmin et al. (2023), the top-ranked method in the DSTC11 competition (Appendix C).
4.2 Robustness Improvement through EPA
The results in Table 2 demonstrate the effectiveness of EPA in enhancing robustness to ASR errors in the DST task. Remarkably, EPA outperformed existing text-based and audio-based augmentation methods across all environments, showing substantial improvement in N-Acc with gains of over 4%, compared to the modest 1-2% improvements of other methods. This trend is particularly pronounced in challenging environments, such as low-accuracy ASR and noisy audio conditions, and is also observed in LLaMA-7B and OPT 6.7B. The ablation study revealed that keyword error augmentation consistently boosted performance across all LLMs, capitalizing on its utility to optimize DST performance against ASR errors.
4.3 Error Analysis
Method Error Type
Wrong Ignore Spurious
Noised Audio
Baseline
$\triangledown$0%
$\triangledown$0%
$\triangledown$0%
EPA w/o Key-aug
$\triangledown$5.29%
$\triangledown$3.72%
$\triangledown$6.31%
EPA
$\triangledown$8.19%
$\triangledown$7.25%
$\triangledown$1.33%
High-acc ASR (WER<0.03)
Baseline
$\triangledown$0%
$\triangledown$0%
$\triangledown$0%
EPA w/o Key-aug
$\triangledown$1.57%
$\triangledown$3.34%
$\triangledown$6.52%
EPA
$\triangledown$7.16%
$\triangledown$2.43%
$\triangledown$0.75%
Table 3: Ablation study with error analysis. Wrong indicates the model predicts incorrect values, Ignore refers to the model ignored mentioned slots, and Spurious denotes the model predicting values for unmentioned slots. Actual error numbers are in parentheses.
To examine the effects of keyword augmentation in more detail, we conducted ablation studies and analyzed error reduction rates by categorizing errors into three types. Table 3 presents the percentage reduction in error rates compared to the baseline. The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%. Interestingly, while keyword augmentation led to substantial reductions in the "Wrong" type error, it increased the "Spurious" error type. This suggests that the model may become overly cautious, occasionally overestimating the presence of named entities.
4.4 Qualitative Assessment of EPA Method
Method Uniq. Words NE.chg [ % ] Audio Sim.[%]
Baseline 1 - -
AEDA Karimi et al. (2021) 1.00$\times$ 44.29 91.57
EDA Wei and Zou (2019) 0.86$\times$ 70.03 61.14
BT Sennrich et al. (2015) 1.21$\times$ 73.46 77.17
TTS-ASR 1.01$\times$ 38.84 98.93
Translating 0.84$\times$ 39.59 94.07
EPA 1.81$\times$ 95.47 91.57
   w/o Keyword Err. 1.57$\times$ 68.81 93.14
Table 4: Assessment of EPA dataset quality across three key aspects (Unique word increased rate(Uniq. Words, Named entity changed rate (NE.chg), and audio similarity (Audio Sim.).
To evaluate the word diversity and audio similarity of the augmented text, we performed a qualitative assessment in Table 4. This assessment includes metrics such as the unique word increase rate, named entity change rate, and audio similarity  [^4] Normalized phoneme edit distance is used for audio similarity calculating.. The results demonstrate that EPA achieves remarkable diversity in unique words (1.81x) and the highest rate of named entity augmentation (95.47%), while maintaining considerable audio similarity (91.57%). Furthermore, keyword augmentation significantly contributes to the improvement, particularly in the named entity change rate (from 68.81% to 91.57%), ensuring that the errors exhibit sufficiently diverse patterns in keywords.
5 Conclusion
We propose a novel data augmentation method tailored for DST tasks that ensures sufficient error patterns in both key phrases and overall text. By leveraging LLMs for their controlled text generation capabilities, we strategically place errors within key phrases. Our method demonstrates substantially improved robustness in DST by generating diverse, plausible keyword errors. Error case analysis reveals that keyword augmentation significantly enhances robustness against ASR errors. As the pioneering research in leveraging LLMs for generating ASR errors, we hope this work lays a strong foundation for future phonetic-based augmentation research.
Limitations
First, the introduction of keyword augmentation, while reducing "Wrong" and "Ignore" errors, resulted in an increase in "Spurious" errors, indicating a potential trade-off in error types. Additionally, our experiments were conducted using only the T5-base model. However, we believe that our method will work effectively with other models as well. Furthermore, we did not explore the application of our method to other similar datasets, such as those used in spoken language understanding (SLU) or named entity recognition (NER) tasks. Future work should consider these areas to validate and extend the applicability of our approach.
Appendix A Baseline Performance Comparison with Clean Text
For comparison, we report the baseline performance on an error-free, clean test dataset. Please note that DSTC11 Soltau et al. (2022) does not provide a text script for the test dataset, so we are manually cleaning 50 dialogues to ensure they are error-free. In the experiment, the baseline model achieved a JGA score of 45.2 % and an N-ACC score of 86.5 % in an ASR error-free environment. Compared to the JGA, which is 34.8 %, and N-ACC, which is 52.07 %, in the ASR-errored environment (High-acc ASR model environment), this discrepancy highlights the significant impact of ASR errors on performance degradation.
Appendix B Training Details
In training models, we used T5-base Roberts et al. (2019) as the backbone model and instructed the model to generate the $B_{t}$ by given $D_{t}$ in sequence to sequence manner, as in Su et al. (2021) and the loss function is
$L=-\sum_{t=1}^{T}\log P(B_{t}|\text{Inst},D_{t}).$ (3)
We set the learning rate as 4e-5 and used the AdamW Loshchilov and Hutter (2017) optimizer. One GeForce RTX 3090 is used for training and the batch size is 16. Trained until reaching the max patient, which is 3.
Appendix C Experiment Details
C.1 Dataset
  • Low-acc ASR environment: Whisper-base model (74M)Radford et al. (2022) is used for transcription. WER on LibriSpeech.test-clean is 0.05.
  • Noisy audio environment: Incorporated authentic cafe and traffic noise from https://freesound.org/ with a 10 to 20 Signal-to-Noise Ratio (SNR) and transcribed it using the Whisper large model.
  • Paraphrased environment: When recording the audio, the text was paraphrased to more closely resemble spoken language Soltau et al. (2022).
  • High-acc ASR environment: Whisper-large model (1550M) is used for transcription. WER on LibriSpeech.test-clean is 0.027.
C.2 Comparison Methods
  • AEDA Karimi et al. (2021): We randomly inserted punctuation marks, effectively maintaining the original word order.
  • EDA Wei and Zou (2019): We augmented data by applying edit-based technique that implements four rule-based modifications—synonym replacement, random insertion, swapping, and deletion.
  • Back Translation Sennrich et al. (2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process. We use English to German 555facebook/wmt19-en-de and German to English666facebook/wmt19-de-en models as translator.
  • TTS-ASR : We used Tacotron2 Shen et al. (2018) for the TTS model to synthesize the audio and use Whisper-base Radford et al. (2022) as an ASR model to simulate the ASR errors.
  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text. Our training set comprised 300 hours of paired clean and ASR-errored text, as detailed in Section 2.2. We fine-tuned the model based on the T5-base architecture Roberts et al. (2019), using the loss function defined in equation 4. The loss function is as follows:
    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (4)
Appendix D Prompt Example
Step 1 Prompt Example
Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
Apply character and word substitutions, additions, or deletions while maintaining the overall pronunciation and context.
Error rate should be high
{Example 1}
Original: they have a single naupliar eye
ASR-errored: they have a single nor pure eye
{Example 2}
Original: i must have saint louis then huzza
ASR-errored: i must have st louis then hazard
{Example 3}
Original: i wonder uncle did not have her come
ASR-errored: i wonder uncle did not have a problem
{Example 4}
Original: i don’t want to have to pay for parking.
ASR-errored:
Step 2 Prompt Example
Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
Here is some example
{Example 1}
Original: I want to buy a book about <hl>luwombo best</hl> restaurant. Keywords : luwombo best
Result: I want to buy a book about luwambo vest restaurant.
{Example 2}
Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>? Keywords : eliot, holiday inn williamsport
Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port
{Example 3}
Original: the <hl>chabuton ramen<hl/> is a restaurant on the east. Keywords : chabuton ramen
Result: the shabuton raymond is a restaurant on the east.
{Example 4}
Original: i was hoping to see local faces while in cambridge. some <hl>entertainment</hl> would grate. Keywords: entertainment
Result:
Method Low-acc ASR Noised Aud. Paraphrased High-acc ASR
JGA N-acc JGA N-acc JGA N-acc JGA N-acc
Baseline 30.05 46.48 29.80 47.49 29.08 48.85 34.73 52.31
AEDA 29.99 46.53 29.80 47.60 28.95 49.02 34.94 52.40
EDA 29.16 47.67 28.93 49.50 28.12 50.10 33.74 53.66
BT 31.54 49.21 31.43 51.25 29.90 51.60 36.25 54.85
TTS-ASR 30.09 46.16 30.32 47.74 29.26 49.31 35.28 51.95
FT 29.98 47.70 29.95 48.49 29.72 50.43 34.82 53.35
EPA (GPT3.5) 32.56 51.62 32.27 53.48 31.10 53.91 36.65 56.14
Baseline 29.82 45.63 29.75 46.20 28.55 48.75 34.83 51.44
AEDA 29.77 46.29 29.69 47.40 29.26 48.93 34.93 52.26
EDA 29.27 47.48 28.57 49.41 28.20 49.99 33.59 53.68
BT 31.66 49.16 31.05 50.81 29.88 51.75 36.19 54.73
TTS-ASR 29.75 45.99 29.80 46.87 29.10 48.12 34.97 51.49
FT 30.86 47.47 30.52 48.53 29.63 50.11 35.73 53.59
EPA (GPT3.5) 32.33 50.74 32.41 52.28 31.02 53.15 36.75 55.82
Baseline 29.77 45.18 29.56 46.61 29.12 48.76 35.05 52.47
AEDA 29.94 46.55 29.72 47.45 29.16 48.62 34.94 52.29
EDA 29.22 47.79 28.59 49.61 27.92 49.88 33.71 53.99
BT 31.86 49.13 31.31 50.87 29.91 51.85 36.38 54.85
TTS-ASR 29.98 46.86 29.84 47.51 28.88 49.20 34.97 52.66
FT 30.37 47.78 29.94 48.32 29.26 50.60 35.20 54.05
EPA (GPT3.5) 32.27 51.01 32.04 52.33 30.72 52.96 36.42 55.66
Table 5: Experiment result with different seeds.

### v2

@@
-Keyword-Aware ASR Error Augmentation for Robust Dialogue State Tracking
+Speak & Spell: LLM-Driven Controllable Phonetic Error Augmentation for Robust Dialogue State Tracking
 Abstract
@@
 As a result, our method generated sufficient error patterns on keywords, leading to improved accuracy in noised and low-accuracy ASR environments.
+[…]
+[…]
+Speak & Spell: LLM-Driven Controllable Phonetic Error Augmentation for Robust Dialogue State Tracking
+Jihyun Lee1, Solee Im3thanks: This work was conducted at POSTECH., Wonjun Lee2, Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence, POSTECH, Republic of Korea 2Department of Computer Science and Engineering, POSTECH, Republic of Korea 3KT {jihyunlee, lee1jun, gblee}@postech.ac.kr {solee.im}@kt.com
 1 Introduction
-Figure 1: Illustration of EPA process.
+Figure 1: Illustration of the EPA process.
+Step 1 introduces overall ASR-like distortions using in-context examples of original–errored pairs, while Step 2 applies keyword-specific phonetic corruptions guided by highlighted slot spans.
+Combining both steps generates realistic and diverse ASR-style errors for robust DST training.
 Task-oriented dialogue systems (TODs) assist users in achieving specific objectives through conversations and are used in various sectors, including customer service and hotel reservations.
@@
 (2013).
-However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR).
+However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR) Pal et al.
+(2020); Kim et al.
+(2021); Yoon et al.
+(2023).
 Notably, Soltau et al.
-(2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments [^1] Joint goal accuracy result of base size model..
+(2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments.
 This decline is primarily due to ASR errors, which frequently misrecognize named entities—a key target in DST Nechaev et al.
@@
 (2021) employed translation model structure to introduce ASR-like errors directly into texts.
-Despite these advancements, prior methods often fail to provide sufficient error patterns for DST model training.
+Huang and Chen (2020) further leveraged word confusion network (WCN) representations for data augmentation to generate acoustically consistent errors.
+Despite these advancements, prior methods often fail to provide sufficient error for DST model training.
 Accurately identifying key terms is vital for DST performance; thus, models need to be trained on a broad spectrum of ASR-errored keywords.
@@
 (2015).
-This oversight results in sub-optimal DST performance against ASR errors.
+This oversight results in sub-optimal performance against ASR errors.
 To address these limitations, we introduce Error Positioning Augmentation (EPA), a straightforward yet effective method that ensures sufficient errors in keywords.
@@
 Despite their strengths, LLMs’ potential for phonetic augmentation remains largely unexplored.
-In our method, we utilize in-context Brown et al.
-(2020) learning to guide the augmentation process.
-By providing examples that highlight the key parts, we enable the model to position errors in keywords.
-Surprisingly, without collecting extensive user voice datasets, a manually created small text dataset (fewer than 10 samples) can generate a wide variety of ASR-errored keywords.
-This simplifies the error generation process.
-Additionally, before keyword-focused augmentation, we augmented the text to incorporate utterance-level overall ASR errors, introducing diverse errors from open-source audio datasets.
-This two-step approach effectively harnesses LLMs for both overall utterance-level and targeted keyword-level phonetic augmentation.
-In our experiments, EPA achieved a diversity in targeted keywords (95.4%) and high variance in overall text (1.81x) while maintaining audio similarity (91.5%).
-During practical DST evaluations, EPA significantly enhanced the robustness of DST models, increasing accuracy from 45.76% to 51.12%.
-Our analysis indicates that these improvements are primarily due to keyword augmentation, which significantly reduces errors in ASR-errored values.
+In our method, we utilize in-context learning Brown et al.
+(2020) with phonetically similar examples to introduce general ASR errors and devise a highlighting method to explicitly localize the error to a target span.
+Surprisingly, without requiring extensive domain-specific user speech data, a publicly available audio dataset and a small set of in-context examples (fewer than 10 samples) are sufficient to generate a wide variety of ASR-errored keywords for DST.
+This significantly simplifies the error generation process.
+In the experiment, to reflect diverse real-world conditions, we evaluated EPA under four ASR environments: a low-accuracy ASR system, noisy audio with café and traffic background, a paraphrased input setting where users naturally rephrased transcriptions, and a high-accuracy ASR system.
+In these experiments, EPA significantly improved model robustness, increasing accuracy from 45.76% to 51.12% with high keyword diversity (95.4%), surpassing the previous best-performing model.
+Our analysis suggests that this improvement is primarily driven by keyword-level augmentation, which effectively mitigates errors in ASR-affected values.
 2 Method
@@
 Before detailing each step, we first clarify the notation.
-Dialogue context from turn 1 to $t$ is denoted as $D_{t}$=$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance.
+Dialogue context from turn 1 to $t$ is denoted as $D_{t}$$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance.
 DST model predicts the dialogue state (also called belief state) $B_{t}$ given $D_{t}$.
@@
 $J$ is the total number of slots.
-Idx Method Examples
-1 Original I’d like to find a vegetarian restaurant, if possible.
-+EPA I’d hike to find a veggie tarian restroom, if possible.
-2 Original I am going to auburn.
-+EPA I am flowing to auburng.
-3 Original Tuesday, going to bailey’s crossroads please.
-+EPA Tuesday, going to baley’s crossroads, peas.
-4 Original Hi! Could you find me a train to loris on thursday?
-+EPA Oh! Could you find me a trai to lorri on thursdae?
-5 Original Ashby is my destination.
-+EPA Ashy’s my desity.
-Table 1: Examples of ASR errors from EPA. We have highlighted utterance level overall errors in yellow and keyword-specific errors in blue.
 2.2 Step 1: ASR Error for Overall Utterance
@@
 We denote this example dataset as $DB=\{(g_{1},e_{1}),\ldots,(g_{I},e_{I})\}$.
-Next, we augment the overall utterance level of $u$ by guiding the LLM using examples retrieved from $DB$ based on audio similarity (see Figure  1, Step 1).
-To calculate audio similarity, we converted the characters to phonemes using the International Phonetic Alphabet (IPA) 222 https://pypi.org/project/eng-to-ipa/ and then computed the similarity between the phonemized $u$ and $DB$.
-After selecting the $k$ examples, we concatenated the instructions, examples, and $u$ into a single prompt and provided it to the LLM.
-This process resulted in the overall ASR-errored user utterance $\dot{u}$.
+Next, we inject errors into $u$ by prompting the LLM with in-context examples.
+We retrieved ($g$, $e$) pairs from the database ($DB$) based on phonetic similarity between $u$ and $g$ (Figure 1, Step 1).
+To compute phonetic similarity, we converted the characters of both $u$ and $g$ into phonemes using the International Phonetic Alphabet (IPA), and calculated similarity using a frequency-based retrieval  [^1] We used BM25Robertson et al.
+(2009), a retrieval model based on term frequency.
+While neural retrievers (e.g., DPRKarpukhin et al.
+(2020)) could be applied, we opted for a frequency-based method, as neural models tend to capture semantic similarity..
+After selecting the top-$k$ ($g$, $e$) pairs, we concatenated the instruction, in-context examples, and $u$ into a single prompt and provided it to the LLM.
+This process results in the overall ASR-errored user utterance, denoted as $\dot{u}$.
 Concretely, $\dot{u}$ can be obtained by
-$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t}).$ (1)
+$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t})$ (1)
+where $\oplus$ denotes concatenation, and we set $k=3$ throughout our experiments. Retrieved examples are provided in Appendix A.3.
 2.3 Step 2: ASR Error for Keywords
+While Step 1 introduces general ASR-style errors into $u$, it does not ensure sufficient error diversity in keyword tokens.
+To construct a more effective training dataset, we explicitly generate keyword-focused ASR errors in Step 2 (Figure 1).
+In this step, we highlight the keywords in $\dot{u}$ using the <hl> tag and instruct the LLM to inject errors specifically within the highlighted spans.
+For the DST task, we treat dialogue state values ($v$) as keywords, although the definition of a keyword may vary depending on the task.
+To facilitate this process, we provide a few examples that illustrate how values within <hl> tags are intended to be modified during augmentation.
+Given these instructions and examples, the LLM generates an augmented utterance $\ddot{u}$ that includes both general and keyword-specific ASR errors.
+Formally, we obtain $\ddot{u}_{t}$ as follows:
+$\ddot{u_{t}}=LLM(Inst_{2}\oplus({g}_{0},{e}_{0})\cdots({g}_{k},{e}_{k})\oplus\dot{u}_{t}).$ (2)
+The used prompts are provided in Appendix A.1.
+2.4 Examples of EPA
+Idx Method Examples
+1 Original Tuesday, going to bailey’s crossroads please.
++EPA Tuesday, going to baley’s crossroads, peas.
+1 Original I’d like to find a vegetarian restaurant, if possible.
++EPA I’d hike to find a veggie tarian restroom, if possible.
+3 Original I am going to auburn.
++EPA I am flowing to auburng.
+4 Original Hi! Could you find me a train to loris on thursday?
++EPA Oh! Could you find me a trai to lorri on thursdae?
+5 Original Ashby is my destination.
++EPA Ashy’s my desity.
+Table 1: Examples of ASR errors from EPA.
+Table 1 shows examples of ASR errors generated by EPA.
+We have highlighted utterance level overall errors in yellow and keyword-specific errors in orange.
+For instance, in Row 1, the model introduces a keyword-level error (bailey’s → baley’s) as well as an additional phonetically plausible insertion (peas), simulating realistic ASR noise.
+Further examples can be found in Appendix 9.
+3 Experiments
 Method Features Low-acc ASR Noised Aud. Paraphrased High-acc ASR
@@
 ASR-translation - - 30.40 47.65 30.14 48.45 29.54 50.38 35.25 53.66
-[…]
+EPA (Opt 6.7B) 31.82 50.73 32.03 51.92 29.57 52.49 37.05 55.78
    w/o Keyword Aug - 31.43 49.63 31.51 50.56 30.41 52.02 36.34 54.57
-EPA (Llama 7B) 31.54 51.12 31.55 52.27 30.10 53.55 36.22 55.49
+EPA (LLaMA2-7B) 31.54 51.12 31.55 52.27 30.10 53.55 36.22 55.49
    w/o Keyword Aug - 31.12 50.33 31.44 52.07 30.01 53.49 35.70 54.90
@@
 Olisia Jacqmin et al. (2023) - - - - 30.17 46.25 30.43 48.07 29.13 49.21 36.1 52.58
-Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments. All results were averaged over three seeds for better consistency.
-While Step 1 introduces general ASR errors into ${u}$, it does not ensure that these errors enhance the diversity of keywords.
-To create a more effectivetraining dataset, we generated keyword-focused ASR errors (see Figure 1, Step 2).Here, we manually created several examples from the original DST dataset by explicitly highlighting the keywords in the clean text using the <hl> tag ($\tilde{g}$) and concatenated them with the ASR-errored versions ($\tilde{e}$).
-We set the $v$ in turn $t$ as keywords.
-We applied the same highlighting to the inference target $\dot{u}$ and then concatenated the instructions, examples and highlighted $\dot{u}$ into a single prompt, as in Step 1.
-We perform the in-context learning with this prompt to generate the augmented utterance $\ddot{u}$, containing overall ASR and keyword-specific errors.
-Explicitly, we get $\ddot{u}$ by
-$\ddot{u_{t}}=LLM(Inst_{2}\oplus(\tilde{g}_{0},\tilde{e}_{0})\cdots(\tilde{g}_{k},\tilde{e}_{k})\oplus\dot{u}_{t}).$ (2)
-We set the example numbers as 3 in both Step1 and Step2.
-After generating the ASR-errored examples, we trained a DST model using both the EPA-generated data and the original seed data.
-Specifically, we fine-tuned T5-base Roberts et al.
-(2019) models for DST tasks, with detailed training procedures outlined in the Appendix B.
-3 Augmentation Results of EPA
-We first assessed the quality of ASR errors simulated by the EPA method to guarantee they closely resembled genuine ASR errors.
-Table 1 demonstrates that with overall utterance and keyword error augmentation, EPA effectively generates sentences that align with our objectives: maintaining phonetic similarity to the original text while confirming errors in keywords.
-4 Experiments
+Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments.
+In the feature columns, “Aud.” indicates the use of audio-based augmentation, “Utt-aug.” denotes utterance-level text augmentation, “Key-aug.” refers to keyword-specific augmentation, and “LLM” indicates whether large language models were used for error generation.
+All results were averaged over three seeds for better consistency.
 3.1 Experimental Setup
@@
 (2019), comprises 8,000 dialogues for training, 1,000 for validation, and 1,000 for testing.
-To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by word error rate (WER) [^3] Based on LibriSpeech.test-clean Radford et al.
-(2022).
-and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model (WER < 0.03) enviroment.
+To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by Word Error Rate (WER) and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model.
 Metrics.
@@
 We also reported named entity accuracy (N-acc), the average accuracy across named entity slots.
-LLMs.
-we primarily used GPT-3.5 Ouyang et al.
-(2022).
-Additionally, as this is the first attempt to augment pronunciation-related errors with LLMs, we also included Llama2-7B Touvron et al.
-(2023) and OPT-6.7B Zhang et al.
-(2022).
 Compared methods.
@@
 Lastly, we included Olisia Jacqmin et al.
-(2023), the top-ranked method in the DSTC11 competition (Appendix C).
+(2023), the top-ranked method in the DSTC11 competition.
+Models.
+For performing EPA, we used diverse types of LLMs, including GPT-3.5Ouyang et al.
+(2022), LLaMA2-7BTouvron et al.
+(2023) and OPT-6.7BZhang et al.
+(2022).
+For the DST task, we fine-tuned a T5-baseRoberts et al.
+(2019) model.
+Further details about the experimental settings are provided in Appendix B.
 3.2 Robustness Improvement through EPA
-The results in Table 2 demonstrate the effectiveness of EPA in enhancing robustness to ASR errors in the DST task.
-Remarkably, EPA outperformed existing text-based and audio-based augmentation methods across all environments, showing substantial improvement in N-Acc with gains of over 4%, compared to the modest 1-2% improvements of other methods.
-This trend is particularly pronounced in challenging environments, such as low-accuracy ASR and noisy audio conditions, and is also observed in LLaMA-7B and OPT 6.7B.
-The ablation study revealed that keyword error augmentation consistently boosted performance across all LLMs, capitalizing on its utility to optimize DST performance against ASR errors.
+EPA improves robustness.
+The results in Table 2 shows the effectiveness of EPA in robustness to ASR errors.
+Remarkably, EPA outperformed existing text-based and audio-based augmentation, showing substantial improvement in JGA and named entity accuracy.
+It also surpassed the previous best-performing model, Olisia, particularly in challenging environments.
+Effectiveness of keyword-specific error.
+In Table 2, we present an ablation study to evaluate the effectiveness of keyword-level augmentation.
+We found that adding keyword-specific ASR errors improved DST performance across all environments and was particularly helpful in enhancing the robustness of named entity accuracy.
+Additional experiments, including generalization to other backbones and tasks, as well as statistical significance analysis, are provided in Appendix C.
+3.3 Comparison with Error Correction Models
+Method Weak-ASR Noised-ASR Strong-ASR
+JGA N-acc JGA N-acc JGA N-acc
+Baseline 29.88 45.76 29.70 46.77 34.87 52.07
++ Correction model 1 23.30 41.50 23.96 42.86 26.68 46.14
++ Correction model 2 30.14 46.79 29.82 47.68 34.62 52.00
++ EPA (with GPT, ours) 32.39 51.12 32.24 52.70 36.61 55.87
+Table 3: Comparison with ASR error correction models under different ASR conditions.
+In Table 3, we compare our method with two ASR error correction approaches.
+For Error Correction Model 1, We fine-tuned a pre-trained on 300 hours of ASR-errored and gold transcription pairs (same as in Section 2.2) in a seq2seq manner, minimizing $L=-\sum_{i=1}^{N}\log P(g_{i}\mid e_{i})$, where $g_{i}$ and $e_{i}$ denote the gold and errored transcriptions, respectively.
+Training used early stopping (patience=3) with batch size 16.
+For Error Correction Model 2, we adopted an off-the-shelf222urlhttps://huggingface.co/oliverguhr/spelling-correction-english-base text correction model to revise typos and transcription errors.
+Each model corrected the test transcriptions from the weak, noised, and strong ASR systems before DST evaluation.
+As in the table, Correction model 1 underperforms the baseline, often altering named entities unnecessarily (e.g., "Grifon" $\rightarrow$ "Bristol"), likely due to limited training diversity.
+Correction model 2 yields minor improvements, while our EPA achieves the highest robustness across all ASR conditions.
+3.4 Qualitative Assessment of EPA Method
+Method
+Uniq.
+Words
+NE.chg
+Pronoun
+Sim.[%]
+Baseline 1 - -
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+[…]
+Table 4: Assessment of EPA dataset quality: Unique word increased rate, Named entity changed rate (NE.chg), and pronunciation similarity.
+Automatic evaluation.
+Although Table 2 confirms EPA’s effectiveness, it remains unclear whether the LLM-generated augmentations truly reflect diverse, keyword-focused ASR-style errors.
+To this end, we perform the quality analysis based on three metrics (Table 4): the unique word increase rate, named entity change rate, and pronunciation similarity with original sentence.
+The results show that EPA achieves remarkable diversity in unique words (1.81×) and the highest named entity change rate (95.47%), while maintaining high pronunciation similarity (91.57%).
+Notably, keyword-level augmentation plays a key role in enhancing named entity variability, increasing the change rate from 68.81% to 95.47%.
+Comparison with Authentic ASR Errors.
+Figure 2: Distribution of edit distance. The x-axis represents edit distances, and the y-axis represents the corresponding ratio.
+Method Text Dist.($\downarrow$) Phoneme Dist.($\downarrow$)
+ASR-L ASR-B ASR-L ASR-B
+TTS-ASR (Whisper-B) 0.030 0.048 0.056 0.061
+TTS-ASR (Whisper-S) 0.030 0.048 0.070 0.077
+ASR trans. (T5-base) 0.025 0.039 0.104 0.116
+EPA (Llama2 7B) 0.218 0.123 0.204 0.256
+EPA (OPT 6.7B) 0.115 0.106 0.071 0.091
+EPA (GPT 3.5) 0.033 0.010 0.009 0.007
+Table 5: Distribution distance (JSD) between Whisper Large/Base model and simulation dataset. (ASR‑L=Whisper Large, ASR‑B=Whisper Base).
+In this analysis, we explored the similarity between simulated ASR errors and authentic ASR errors from the perspective of edit distance.
+Specifically, we examined the distribution of edit distances in simulated data and in errors produced by Whisper large/base models, considering both character- and phoneme-level representations (Figure 2).
+To quantify the distributional differences, we computed the Jensen-Shannon Divergence (JSD), a symmetric variant of the Kullback-Leibler divergence (Table 5).
+Our experiments yielded several noteworthy findings.
+Notably, the LLM-simulated errors closely matched the distribution of real ASR errors, particularly at the phoneme level.
+This suggests that LLMs can effectively capture pronunciation-level variations and generate realistic ASR-style distortions.
+In contrast, errors produced by TTS-based synthesis or ASR-translation models exhibited larger divergence from real ASR patterns, likely due to their tendency to generate more diverse but less phonetically grounded outputs.
+Human evaluation.
+To further verify the quality of our EPA method, we conducted a human evaluation using 100 sentence pairs, each consisting of an original sentence and its augmented counterpart, with two human evaluators.
+Participants rated how likely the change resembled an ASR error on a 4-point Likert scale, where 1 indicated "not like an ASR error" and 4 indicated "clearly an ASR error." The average rating was 3.22 with moderate inter-rater agreement (Gwet’s AC2 Gwet (2008) = 0.590), suggesting that most EPA-generated edits were perceived as realistic ASR errors.
+Details on the evaluation metric and human evaluation are provided in Appendix D.
 3.5 Error Analysis
+We additionally analyze the impact of keyword augmentation by examining how it influences specific error types in DST predictions.
+Table 6 presents the percentage reduction in error rates compared to the baseline.
+The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%.
+Interestingly, while keyword augmentation led to substantial reductions in "Wrong" errors, it also caused a slight increase in "Spurious" errors.
+This may be because the model, after repeatedly seeing phonetic noise around slot values, becomes overly sensitive and starts hallucinating unmentioned slots.
+A potential mitigation is to introduce an additional loss term for slot presence prediction Heck et al.
+(2020); Kim et al.
+(2019) , helping the model better distinguish between mentioned and unmentioned slots.
 Method Error Type
@@
 $\triangledown$1.33%
-High-acc ASR (WER<0.03)
-Baseline
-$\triangledown$0%
-$\triangledown$0%
-$\triangledown$0%
-EPA w/o Key-aug
-$\triangledown$1.57%
-$\triangledown$3.34%
-$\triangledown$6.52%
-EPA
-$\triangledown$7.16%
-$\triangledown$2.43%
-$\triangledown$0.75%
 Table 6: Ablation study with error analysis.
-Wrong indicates the model predicts incorrect values, Ignore refers to the model ignored mentioned slots, and Spurious denotes the model predicting values for unmentioned slots.
+Wrong indicates the model predicts incorrect values, Ignore refers to ignored mentioned slots, and Spurious denotes predicting values for unmentioned slots.
 Actual error numbers are in parentheses.
-To examine the effects of keyword augmentation in more detail, we conducted ablation studies and analyzed error reduction rates by categorizing errors into three types.
-Table 3 presents the percentage reduction in error rates compared to the baseline.
-The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%.
-Interestingly, while keyword augmentation led to substantial reductions in the "Wrong" type error, it increased the "Spurious" error type.
-This suggests that the model may become overly cautious, occasionally overestimating the presence of named entities.
-4.4 Qualitative Assessment of EPA Method
-Method Uniq. Words NE.chg [ % ] Audio Sim.[%]
-Baseline 1 - -
-AEDA Karimi et al. (2021) 1.00$\times$ 44.29 91.57
-EDA Wei and Zou (2019) 0.86$\times$ 70.03 61.14
-BT Sennrich et al. (2015) 1.21$\times$ 73.46 77.17
-TTS-ASR 1.01$\times$ 38.84 98.93
-Translating 0.84$\times$ 39.59 94.07
-EPA 1.81$\times$ 95.47 91.57
-   w/o Keyword Err. 1.57$\times$ 68.81 93.14
-Table 4: Assessment of EPA dataset quality across three key aspects (Unique word increased rate(Uniq. Words, Named entity changed rate (NE.chg), and audio similarity (Audio Sim.).
-To evaluate the word diversity and audio similarity of the augmented text, we performed a qualitative assessment in Table 4.
-This assessment includes metrics such as the unique word increase rate, named entity change rate, and audio similarity  [^4] Normalized phoneme edit distance is used for audio similarity calculating..
-The results demonstrate that EPA achieves remarkable diversity in unique words (1.81x) and the highest rate of named entity augmentation (95.47%), while maintaining considerable audio similarity (91.57%).
-Furthermore, keyword augmentation significantly contributes to the improvement, particularly in the named entity change rate (from 68.81% to 91.57%), ensuring that the errors exhibit sufficiently diverse patterns in keywords.
 4 Conclusion
@@
 Limitations
-First, the introduction of keyword augmentation, while reducing "Wrong" and "Ignore" errors, resulted in an increase in "Spurious" errors, indicating a potential trade-off in error types.
-Additionally, our experiments were conducted using only the T5-base model.
-However, we believe that our method will work effectively with other models as well.
-Furthermore, we did not explore the application of our method to other similar datasets, such as those used in spoken language understanding (SLU) or named entity recognition (NER) tasks.
-Future work should consider these areas to validate and extend the applicability of our approach.
-Appendix A Baseline Performance Comparison with Clean Text
-For comparison, we report the baseline performance on an error-free, clean test dataset.
-Please note that DSTC11 Soltau et al.
-(2022) does not provide a text script for the test dataset, so we are manually cleaning 50 dialogues to ensure they are error-free.
-In the experiment, the baseline model achieved a JGA score of 45.2 % and an N-ACC score of 86.5 % in an ASR error-free environment.
-Compared to the JGA, which is 34.8 %, and N-ACC, which is 52.07 %, in the ASR-errored environment (High-acc ASR model environment), this discrepancy highlights the significant impact of ASR errors on performance degradation.
-Appendix B Training Details
+Through detailed error analysis, we identified a trade-off introduced by our keyword-focused phonetic augmentation strategy.
+While the augmentation helps the model become more robust to noisy slot expressions—leading to substantial reductions in "Wrong" errors—it also increases the model’s sensitivity to phonetic variations.
+As a result, we observed cases where the model hallucinates slot values that were not actually mentioned, thereby increasing the number of "Spurious" errors.
+This hallucination effect represents a key limitation of our method.
+We attribute it to the model’s repeated exposure to noisy keywords, which may cause it to overgeneralize phonetic cues as valid slot mentions.
+As a direction for future work, we plan to incorporate an auxiliary loss term for slot presence prediction Heck et al.
+(2020); Kim et al.
+(2019) to help the model better distinguish between mentioned and unmentioned slots and mitigate this side effect.
+Ethical Considerations
+Our phonetic augmentation method, while effective for simulating ASR-style errors, may raise several ethical concerns.
+One such concern is the potential for accent bias, wherein phonetic transformations may disproportionately reflect majority or standard pronunciations, thereby marginalizing regional or minority accents.
+Another concern is the inadvertent corruption of proper names, particularly those that are less common or culturally specific, which could lead to misrepresentation or reduced inclusivity.
+We acknowledge these risks and emphasize that our method relies on LLMs trained on diverse and large-scale corpora.
+As such, the phonetic errors generated are likely to reflect dominant patterns present in mainstream ASR systems, rather than rare or region-specific variations.
+Nonetheless, we recognize the importance of fairness and inclusivity in language technologies and believe that future work should explore augmentation strategies that are more sensitive to accent and cultural variability.
+Acknowledgements
+This research was supported by the Culture, Sports and Tourism R&D Program through the Korea Creative Content Agency funded by the Ministry of Culture, Sports and Tourism in 2025 (Project No.
+RS-2025-02413038, 45%), the IITP-ITRC (Information Technology Research Center) Program funded by the Ministry of Science and ICT (Project No.
+IITP-2025-RS-2024-00437866, 45%), and the IITP Artificial Intelligence Graduate School Program (POSTECH) funded by the Ministry of Science and ICT (Project No.
+RS-2019-II191906, 10%).
+Appendix A Details of the EPA Method
+A.1 Prompt Used for EPA
+The prompts used in Step 1 and Step 2 are provided below.
+Step 1 Prompt Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
+Apply character and word substitutions, additions, or deletions while maintaining the overall pronunciation and context.
+Error rate should be high
+. Example 1
+Original: they have a single naupliar eye
+ASR-errored: they have a single nor pure eye Example 2
+Original: i must have saint louis then huzza
+ASR-errored: i must have st louis then hazard Example 3
+Original: i wonder uncle did not have her come
+ASR-errored: i wonder uncle did not have a problem
+. Now, following the above examples, generate an ASR-errored version of the following sentence:
+Original: [Target utterance]
+ASR-errored:
+Step 2 Prompt Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
+Here is some example.
+. Example 1
+Original: I want to buy a book about <hl>luwombo best</hl> restaurant.
+Keywords : luwombo best Result: I want to buy a book about luwambo vest restaurant. Example 2
+Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>?
+Keywords : eliot, holiday inn williamsport Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port Example 3
+Original: the <hl>chabuton ramen<hl/> is a restaurant on the east.
+Keywords : chabuton ramen Result: the shabuton raymond is a restaurant on the east.
+. Now, following the above examples, generate an ASR-errored version of the following sentence:
+Original: [Target utterance with <hl> tag]
+ASR-errored:
+A.2 Detailed Keyword Highlighting Strategy for EPA
+Example of adding <hl> tag
+Original Hi, I need to go to Green Day hotel, then book a table at the Grill House.
+Dialogue State hotel-name: Green Day, restaurant-name:Grill House
+With <hl> tags Hi, I need to go to <hl>Green Day</hl> hotel, then book a table at the <hl>Grill House</hl>.
+Table 7: Example of keyword highlighting using <hl> tags based on dialogue state annotations.
+To explicitly introduce keyword-specific ASR errors, we first identify dialogue state values from the training corpus and match them against the user utterance ($\dot{u}$).
+Matched values are then automatically wrapped with <hl> tags based on slot annotations (e.g., DST slot labels or NER tags), as shown in Table 7.
+These highlighted utterances are passed to the LLM, which is instructed to perturb the text within the <hl> tags while preserving the rest.
+This keyword highlighting strategy is task-agnostic and can be easily applied to other keyword-sensitive tasks such as Named Entity Recognition (NER) or Spoken Language Understanding (SLU), where certain slot values or entities are critical for downstream prediction.
+Example 1
+Target can you tell me the address to the police station in point pleasant?
+Retrieved 1 frayser station was not the depot on the point
+ + ASR freya station was not the deep watch on the point
+Retrieved 2 can you get me the maldeamores saga
+ + ASR can you get me the melamorphos
+Retrieved 3 cannot you tell her whom i am eh joseph
+ + ASR cannot you tell her whom i am
+Example 2
+Target no, i just need to make sure it’s cheap. oh, and i need parking.
+Retrieved 1 i need fifty ten foot long segments of wire
+ + ASR i need fifty ten foot long signals of my life
+Retrieved 2 a drive with a different encoding mechanism would need different patterns
+ + ASR and drive was a different building recognition would need different patterns
+Retrieved 3 to reach to calcutta you need less time to reach dhaka
+ + ASR to reach tukaukara you need last time to reach daka
+Example 3
+Target i’m open to any kind of food. i’m looking for something in the centre and on the expensive side.
+Retrieved 1 kokai means open to the public or laid open
+ + ASR cook eye means open to the public all laid open
+Retrieved 2 the town of beauharnois was the major centre
+ + ASR the town of bo hanwa was the major center
+Retrieved 3 the gate is open at eleven
+ + ASR the gate is open at 11
+Table 8: In-context examples retrieved based on phoneme-level similarity.
+For each target utterance (top row), we retrieve three ($g$, $e$) example pairs from the database using phonetic similarity between the target and $g$.
+Colored segments highlight phonetically similar phrases between the target and retrieved examples.
+A.3 Retrieved In-Context Example
+In Section 2.2, we retrieved in-context examples based on phoneme-level similarity.
+Table 8, we present several representative examples to illustrate this retrieval process, showing how phonetically similar phrases (highlighted in color) are matched between the target and retrieved utterances.
+This demonstrates that the retrieval mechanism effectively captures pronunciation-level patterns relevant to ASR-style errors.
+A.4 Additional Examples of ASR-style Errors
+Original Augmented
+- no, i just need to make sure it’s cheap. oh, and i need parking - no, i just need to make sure it’s sheep. oh, and i need parking.
+- i am departing from marion - - i am departing from maryland
+- no, i don’t need anything else right now. thank you for your assistance. good bye. - no, i don’t need anyone else right now. thank you for your persistence. good buy.
+- i would like to go to sandy please. - i would like to go to cindy please.
+- i would like to keep it in the moderate range, please. - i would like to keep it in the mod rain, please
+- could i get the address for it? i would also like an expensive place to eat around it. - could i get the actress for it? i would also like an extensive place to eat around it.
+- i need to take a train out of garrett, i will be leaving town on wednesday - i need to make a plane out of garrett, i will be weaving town on wednesday.
+- do you have any indian restaurants in the south in a different price range? - do you have any indonesian restaurants in the south in a different prize range
+- nope, same people. - nope, same pupil.
+- i’m looking for a college type attraction - i’m looking for a knowledge-type action.
+- yes, please book me a room for friday - yes, please cook me a broom for friday
+- yes, could you please email or fax me the fare amount, as well as the reference number? - yes, could you please email or text me the fair amount, as well as thereference code?
+- ois el shaddai a guest house or hotel? iz let shadai a gest house or motel?
+- great! i also need a train from mount pleasant to sabattus, please. - great! i also need a strain from mount pleasant to suspicious, please
+- yes, can you help me find a train that can take me from lovelock to abbot? - yes, can you help me find a plane that can take me from love lock to rabbit?
+Table 9: Examples of augmented utterances generated by injecting phoneme-level ASR-style errors.
+For each original utterance (left), the corresponding augmented version (right) includes substitutions that mimic realistic ASR recognition mistakes.
+Blue-colored phrases indicate changes in keywords that are used as slot values in DST, while orange-colored phrases represent overall ASR-style errors.
+Table 9 presents additional examples of ASR-style errors generated by our EPA method, including both general and keyword-specific transformations.
+Appendix B Experimental Setup
+B.1 Details of the ASR Environment
+  • Low-acc ASR environment: Whisper-base model (74M)Radford et al. (2022) is used for transcription. WER on LibriSpeech.test-clean is 0.05.
+  • Noisy audio environment: Incorporated authentic cafe and traffic noise from https://freesound.org/ with a 10 to 20 Signal-to-Noise Ratio (SNR) and transcribed it using the Whisper large model.
+  • Paraphrased environment: When recording the audio, the text was paraphrased to resemble more natural, real-life spoken languageSoltau et al. (2022).
+  • High-acc ASR environment: Whisper-large model (1550M) is used for transcription. WER on LibriSpeech.test-clean is 0.027.
+B.2 Comparison Methods
+  • AEDA Karimi et al. (2021): We randomly inserted punctuation marks, effectively maintaining the original word order.
+  • EDA Wei and Zou (2019): We augmented data by applying edit-based technique that implements four rule-based modifications—synonym replacement, random insertion, swapping, and deletion.
+  • Back Translation Sennrich et al.
+(2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process.
+We use English to German 333facebook/wmt19-en-de and German to English444facebook/wmt19-de-en models as translator.
+  • TTS-ASR : We used Tacotron2 Shen et al. (2018) for the TTS model to synthesize the audio and use Whisper-base Radford et al. (2022) as an ASR model to simulate the ASR errors.
+  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text.
+Our training set comprised 300 hours of paired clean and ASR-errored text.
+We fine-tuned the model based on the T5-base architecture Roberts et al.
+(2019), using the loss function defined in equation 3.
+The loss function is as follows:
+    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (3)
+B.3 Training Details
 In training models, we used T5-base Roberts et al.
@@
 Trained until reaching the max patient, which is 3.
-Appendix C Experiment Details
-C.1 Dataset
-  • Low-acc ASR environment: Whisper-base model (74M)Radford et al. (2022) is used for transcription. WER on LibriSpeech.test-clean is 0.05.
-  • Noisy audio environment: Incorporated authentic cafe and traffic noise from https://freesound.org/ with a 10 to 20 Signal-to-Noise Ratio (SNR) and transcribed it using the Whisper large model.
-  • Paraphrased environment: When recording the audio, the text was paraphrased to more closely resemble spoken language Soltau et al. (2022).
-  • High-acc ASR environment: Whisper-large model (1550M) is used for transcription. WER on LibriSpeech.test-clean is 0.027.
-C.2 Comparison Methods
-  • AEDA Karimi et al. (2021): We randomly inserted punctuation marks, effectively maintaining the original word order.
-  • EDA Wei and Zou (2019): We augmented data by applying edit-based technique that implements four rule-based modifications—synonym replacement, random insertion, swapping, and deletion.
-  • Back Translation Sennrich et al.
-(2015): We translated original texts to error texts and then back to the original texts for generating syntactic variations during the process.
-We use English to German 555facebook/wmt19-en-de and German to English666facebook/wmt19-de-en models as translator.
-  • TTS-ASR : We used Tacotron2 Shen et al. (2018) for the TTS model to synthesize the audio and use Whisper-base Radford et al. (2022) as an ASR model to simulate the ASR errors.
-  • ASR translation: We employed a sequence-to-sequence structure to translate clean text into ASR-errored text.
-Our training set comprised 300 hours of paired clean and ASR-errored text, as detailed in Section 2.2.
-We fine-tuned the model based on the T5-base architecture Roberts et al.
-(2019), using the loss function defined in equation 4.
-The loss function is as follows:
-    $L=-\sum_{i=1}^{I}\log P(e_{i}|g_{i}).$ (4)
-Appendix D Prompt Example
-Step 1 Prompt Example
-Generate ASR error augmented text with similar pronunciation but different words based on the given gold text examples.
-Apply character and word substitutions, additions, or deletions while maintaining the overall pronunciation and context.
-Error rate should be high
-{Example 1}
-Original: they have a single naupliar eye
-ASR-errored: they have a single nor pure eye
-{Example 2}
-Original: i must have saint louis then huzza
-ASR-errored: i must have st louis then hazard
-{Example 3}
-Original: i wonder uncle did not have her come
-ASR-errored: i wonder uncle did not have a problem
-{Example 4}
-Original: i don’t want to have to pay for parking.
-ASR-errored:
-Step 2 Prompt Example
-Change the key words in <hl> tag, to having a ASR error. ASR error has similar pronounciation with the correct word, but different charater.
-Here is some example
-{Example 1}
-Original: I want to buy a book about <hl>luwombo best</hl> restaurant. Keywords : luwombo best
-Result: I want to buy a book about luwambo vest restaurant.
-{Example 2}
-Original: hi, i’m looking for a bus that is depart from <hl>eliot<hl/> and arriving to <hl>holiday inn williamsport<hl/>? Keywords : eliot, holiday inn williamsport
-Result: hi, i’m looking for a bus that is depart from Ellyot and arriving to holliday inn william’s port
-{Example 3}
-Original: the <hl>chabuton ramen<hl/> is a restaurant on the east. Keywords : chabuton ramen
-Result: the shabuton raymond is a restaurant on the east.
-{Example 4}
-Original: i was hoping to see local faces while in cambridge. some <hl>entertainment</hl> would grate. Keywords: entertainment
-Result:
+Appendix C Further Experiments
+C.1 Statistical Significance Analysis
+Method Low-acc ASR Noised Aud. Paraphrased High-acc ASR
+JGA N-acc JGA N-acc JGA N-acc JGA N-acc
+Baseline
+AEDA ns ns ns ns ns ns ns ns
+EDA * * ** ** ns ** ** **
+BT ** ** ** ** * ** ** *
+TTS-ASR ns ns ns ns ns ns ns ns
+ASR trans. ns * ns ns ns * * *
+EPA (GPT-3.5) *** ** *** *** ** ** * **
+Table 10: Statistical significance results compared to the Baseline using paired t-tests across three random seeds.
+Stars indicate significance levels: * for $p<0.05$, ** for $p<0.01$, *** for $p<0.001$, and ns for non-significant differences.
+To assess the reliability of our results, we conducted paired t-tests between each method and the Baseline to determine whether the observed performance improvements are statistically significant.
+We report 95% confidence intervals to reflect performance variability.
+Statistical significance is denoted using asterisks: $\ast$ for $p<0.05$, $\ast\ast$ for $p<0.01$, and $\ast\ast\ast$ for $p<0.001$.
+As shown in Table 10, EPA (GPT-3.5) achieves statistically significant gains in nearly all evaluation settings, particularly under low-accuracy and noised ASR conditions.
+These results confirm that the improvements brought by our method are both consistent and statistically reliable.
+C.2 Baseline Performance Comparison with Clean Text
+For comparison, we report the baseline performance on an error-free, clean test dataset.
+Please note that DSTC11 Soltau et al.
+(2022) does not provide a text script for the test dataset, so we are manually cleaning 50 dialogues to ensure they are error-free.
+In the experiment, the baseline model achieved a JGA score of 45.2 % and an N-ACC score of 86.5 % in an ASR error-free environment.
+Compared to the JGA, which is 34.8 %, and N-ACC, which is 52.07 %, in the ASR-errored environment (High-acc ASR model environment), this discrepancy highlights the significant impact of ASR errors on performance degradation.
+C.3 Effect of Keyword Augmentation
+Method Weak-ASR Noised Audio Strong ASR
+JGA N-acc JGA N-acc JGA N-acc
+Baseline 29.88 45.76 29.70 46.77 34.87 52.07
++ Overall Error (Step 1) 31.31 50.67 31.13 52.29 35.40 55.80
++ Keyword Error (Step 2) 31.66 49.88 31.61 51.02 36.12 54.23
+EPA (Step1+2) 32.39 51.12 32.24 52.70 36.61 55.87
+Table 11: Ablation on overall and keyword-specific augmentation steps. Combining both (EPA) yields the highest robustness across ASR conditions.
+To further validate the contribution of each component in our method, we conducted an ablation study separating Step 1 (overall error augmentation) and Step 2 (keyword-specific error augmentation), using GPT-3.5 as the generator.
+Both augmentation strategies individually improved performance over the baseline.
+Step 1 yielded consistent gains across all ASR settings, while Step 2 notably enhanced slot-value accuracy.
+When combined (EPA, ours), the model achieved the highest robustness in all evaluation conditions, confirming that the two strategies are complementary.
+C.4 Experiments with a Different Baseline
+Method Low-acc ASR Noised Aud. Paraphrased
+JGA N-acc JGA N-acc JGA N-acc
+Baseline 29.9 45.82 27.25 41.81 25.81 44.51
+   + EPA 30.63 48.54 27.5 46.52 27.65 47.96
+Table 12: Experiment with GPT-2 model as baseline.
+In the main experiments, we use T5-base as the backbone model.
+To assess the generalizability of our approach, we additionally conduct experiments using a GPT-2 Radford et al.
+(2019) model, as shown in Table 12.
+The results show a consistent trend with those of T5-base, demonstrating that our method is effective across different backbone architectures.
+C.5 Generalizability to Other Tasks
+Method Low-acc ASR Noised Audio High-acc ASR
+Baseline 56.29 60.50 60.02
++ OPT 6.7B 59.64 62.84 62.47
++ LLaMA 7B 58.29 60.53 60.46
++ GPT-3.5 (125B) 59.39 62.62 62.16
+Table 13: NER results on the ASAPP/SLUE dataset under different ASR conditions. We reported the F1 score.
+Method Low-acc ASR Noised Audio High-acc ASR
+Baseline 55.25 63.23 64.32
++ OPT 6.7B 56.97 64.94 66.10
++ LLaMA 7B 57.24 64.26 65.56
++ GPT-3.5 (125B) 58.78 66.16 67.49
+Table 14: SLU results on the SLURP dataset under different ASR conditions. We reported F1 score.
+To evaluate the generalizability of our approach beyond the DST domain, we extended our experiments to two additional spoken language understanding tasks: Named Entity Recognition (NER) and Spoken Language Understanding (SLU).
+We applied our EPA methodology under three ASR conditions—low-accuracy ASR, noised audio, and high-accuracy ASR—using the same experimental setup as in the DSTC11 evaluation.
+We used asapp/slue dataset for NER taskShon et al.
+(2022), and SLURP dataset Bastianelli et al.
+(2020) for SLU task.
+The results, shown in Table 13 and Table 14, demonstrate that our method consistently improves performance across all ASR conditions for both NER and SLU tasks.
+Notably, the gains are especially prominent under low-accuracy and noisy conditions, confirming that our approach is broadly applicable to other tasks.
+C.6 Additional Fine-grained Metrics
+Method Precision Recall F1 Slot Accuracy
+Baseline 50.5 50.1 50.3 92.8
++ TTS-ASR 49.5 49.6 49.5 93.4
++ ASR-Translation 51.8 51.7 51.7 93.1
++ EPA (GPT3.5) 55.2 54.8 55.0 93.7
+Table 15: Fine-grained DST metrics under Low-accuracy ASR setting.
+Method Precision Recall F1 Slot Accuracy
+Baseline 52.0 51.3 51.6 92.5
++ TTS-ASR 53.7 53.1 53.4 92.8
++ ASR-Translation 53.8 53.0 53.4 92.5
++ EPA (GPT3.5) 56.1 55.3 55.7 93.2
+Table 16: Fine-grained DST metrics under Noised ASR setting.
+To supplement the main results focusing on JGA and N-Acc, we report additional fine-grained metrics—Precision, Recall, F1, and Slot Accuracy—under two ASR corruption settings: Low-accuracy ASR and Noised ASR.
+These metrics provide a more comprehensive view of model behavior in diverse error conditions in table 15 and  16.
+C.7 Results with Different Random Seeds
 Method Low-acc ASR Noised Aud. Paraphrased High-acc ASR
@@
 TTS-ASR 30.09 46.16 30.32 47.74 29.26 49.31 35.28 51.95
-FT 29.98 47.70 29.95 48.49 29.72 50.43 34.82 53.35
+ASR trans. 29.98 47.70 29.95 48.49 29.72 50.43 34.82 53.35
 EPA (GPT3.5) 32.56 51.62 32.27 53.48 31.10 53.91 36.65 56.14
@@
 TTS-ASR 29.75 45.99 29.80 46.87 29.10 48.12 34.97 51.49
-FT 30.86 47.47 30.52 48.53 29.63 50.11 35.73 53.59
+ASR trans. 30.86 47.47 30.52 48.53 29.63 50.11 35.73 53.59
 EPA (GPT3.5) 32.33 50.74 32.41 52.28 31.02 53.15 36.75 55.82
@@
 TTS-ASR 29.98 46.86 29.84 47.51 28.88 49.20 34.97 52.66
-FT 30.37 47.78 29.94 48.32 29.26 50.60 35.20 54.05
+ASR trans. 30.37 47.78 29.94 48.32 29.26 50.60 35.20 54.05
 EPA (GPT3.5) 32.27 51.01 32.04 52.33 30.72 52.96 36.42 55.66
 Table 17: Experiment result with different seeds.
+Table 17 reports the results of our main experiments (Table 2) repeated with three different random seeds, demonstrating the consistency of the observed trends.
+Appendix D Details of Quality Evaluation
+D.1 About Metric
+As described in Section 3.4, we use a phonetic similarity metric to evaluate pronunciation-level consistency between the original and augmented text.
+Specifically, we compute the normalized phoneme edit distance, which quantifies the minimal number of phoneme-level operations required to transform one utterance into another.
+A higher score indicates greater phonetic similarity.
+We used the eng-to-ipa library [^5] https://pypi.org/project/eng-to-ipa/ for phoneme conversion in our implementation, as shown in the code snippet below.
+def phonetic_similarity(original_text, augmented_text):
+original_ipa = to_phoneme(original_text)
+augmented_ipa = to_phoneme(augmented_text)
+edit_distance = nltk.edit_distance(original_ipa, augmented_ipa)
+# Normalize the edit distance
+max_length = max(len(original_ipa), len(augmented_ipa))
+normalized_distance = float(edit_distance) / float(max_length)
+# Convert to similarity score
+similarity_score = 1 - normalized_distance
+return similarity_score
+D.2 Human Evaluation Details
+To assess the plausibility of the generated ASR-style errors, we conducted a human evaluation involving two graduate students.
+Each participant was asked to rate whether a given sentence transformation could plausibly be attributed to an ASR error, using a 4-point Likert scale:
+  • 1 – Not at all: The change is unlikely to be due to an ASR error. It appears to stem from other factors such as meaning alteration or stylistic variation.
+  • 2 – Unlikely: The transformation is probably not caused by an ASR error.
+  • 3 – Somewhat likely: The transformation may plausibly be caused by an ASR error.
+  • 4 – Very likely: The transformation clearly appears to result from an ASR error.
+Each sentence pair (original and transformed) was rated independently by both annotators.
+Inter-rater agreement and average scores are reported in Section 3.4.
+The distribution of Likert scores for each annotator is as follows : Annotator 1 assigned 5% of scores as 1, 7% as 2, 17% as 3, and 71% as 4.
+Annotator 2 assigned 3% of scores as 1, 24% as 2, 53% as 3, and 20% as 4.